The breakdown voltage of a randomly chosen diode of a certain type is known to be normally distributed with mean value 39 V and standard

deviation 1.1 V.

a) What is the probability that the voltage of a single diode is between 37.98 and 41.24?

b) What value c is such that only 1% of all diodes have voltages exceeding that value?

c) If 6 diodes are independently selected, what is the probability that at least 2 has a voltage exceeding 41.24?


1
Expert's answer
2021-12-20T16:15:57-0500

Sol) Given,

Mean =39, S. D=1.1

a)

"\begin{aligned} & P(37.98<x<41.24) \\ & P(x<42)-P(x<39) \\ z=& \frac{x-\text { mean }}{\text { S.D }} \\=& P\left(z<\frac{42-40}{1.5}\right)-P\left(z<\frac{39-40}{1.5}\right) \\=& P(z<1.33333)-P(z<-0.66666) \\=& 0.9088-0.2525=0.6563 \end{aligned}"

 

b) 1% of diodes

"z=\frac{x-\text { mean }}{S \cdot D}\Rightarrow \quad 0.1=\frac{x-\text { mean }}{S \cdot D}"

z-value corresponding to 0.15 is 1.04

"\begin{aligned}

1.04 &=\frac{x-40}{15} \\

(1.04)(15) & \neq 40=x \\

x &=41.56

\end{aligned}"

 

c)

"\begin{aligned}
P \text { (atleast } 41.24 \text { ) } &=1-P(\text { less then } 41.24) \\
&=1-P(x<41.24) \\
&=1-P\left(z<\frac{41.24-40}{1.5}\right) \\
&=1-P(z<.826) \\
&=1-0.5000 \\
&=0.5 \\
\text{As 6 diode are selected}\\
&=1-(0.5)^{6} \\
&=1-0.015625 \\
&=0.9843
\end{aligned}"